OpenStudy (anonymous):

Find the integral of 1/((x+1)(x^2+1)) dx.

OpenStudy (kc_kennylau):

Have you learnt partial fractions?

OpenStudy (anonymous):

Yes, but this one is weird.

OpenStudy (kc_kennylau):

Why?

OpenStudy (anonymous):

Because x^2+1 doesn't have real roots.

OpenStudy (kc_kennylau):

Then treat it as one entity

OpenStudy (anonymous):

resolve into partial fractions and integate

OpenStudy (anonymous):

can you show the partial fractions?

OpenStudy (anonymous):

\[\frac{ 1 }{ \left( x+1 \right)\left( x^2+1 \right) }=\frac{ A }{ x+1 }+\frac{ Bx+C }{ x^2+1 }\] can you proceed further?

OpenStudy (kc_kennylau):

This result would be useful: http://www.wolframalpha.com/input/?t=crmtb01&f=ob&i=d%2Fdx%20tan%5E-1(x)

OpenStudy (anonymous):

\[1=A(x^2+1)+(Bx+C)(x+1)\] \[1=A(x^2+1)+Bx^2+Bx+Cx+C\] Equate the coefficients of x^2,x and constant terms

OpenStudy (anonymous):

0=A+B 0=B+C 1=A+C find A,B,C and then integrate.